OpenStudy (anonymous):

Find the equation of the tangent line to the graph f(x)=(x)/square root of (7-3x) at the point where x=1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

First differentiate the equation to find y' the use slope point form (y-y1)/(x-x1)=y'

OpenStudy (anonymous):

how do you take off the square root

OpenStudy (anonymous):

why do you need to take off the square root ?differential of sqrt(x)=1/2sqrt(x)

OpenStudy (anonymous):

see the differntial of the equation is y'=(sqrt(7-3x)+(3x/sqrt(7-3x)))/(7-3x) now put x=1 y'=(2+(3/2))/2 so then slope = 7/4

OpenStudy (anonymous):

ohhhhh I GET IT thx :)

OpenStudy (anonymous):

ohh but im supose to give the equation of the tangent line ( y=mx+b) if I have only the slope and how am i supose to find the b

OpenStudy (anonymous):

now you know slope of the line then , x=1 so y=1/sqrt(4),y=1/2 so now you have a point and you have the slope.